OpenStudy (anonymous):

Suppose 5 cards are drawn from a standard 52-card deck. What is the probability of drawing four Aces?

OpenStudy (dumbcow):

prob = num of aces/num of cards = 4/52 * 3/51 * 2/50 * 1/49 = 1/13 * 1/17 * 1/25 *1/49